In a machine shop, a technician cuts a 2-meter by 3-meter rectangular sheet of metal into two equal triangular pieces. What is the diagonal measurement of each triangle? Can these pieces be slightly modified so the diagonals are exactly 2.5 meters?


1
Expert's answer
2021-03-26T13:48:31-0400

"d=\sqrt{2^2+3^2}=\sqrt{13}=3.61\;m."

If we cut the sheet into 2 equal rectangles 2-meter by 1.5-meter and then cut each two equal triangular pieces, the diagonal measurement of each triangle will be

"d=\sqrt{2^2+1.5^2}=\sqrt{6.25}=2.5\;m."